### v1

@@ Abstract
A low-rank approximation of a parameter-dependent matrix $A(t)$ is an important task in the computational sciences appearing for example in dynamical systems and compression of a series of images. In this work, we introduce AdaCUR, an efficient algorithm for computing a low-rank approximation of parameter-dependent matrices via CUR decomposition. The key idea for this algorithm is that for nearby parameter values, the column and row indices for the CUR decomposition can often be reused. AdaCUR is rank-adaptive, certifiable and has complexity that compares favourably against existing methods. A faster algorithm which we call FastAdaCUR that prioritizes speed over accuracy is also given, which is rank-adaptive and has complexity which is at most linear in the number of rows or columns, but without certification.
@@ 1 Introduction
The task of finding a low-rank approximation to a matrix is ubiquitous in the computational sciences [56]. In large scale problems, low-rank approximation provides an efficient way to store and process the matrix. In this work, we study the low-rank approximation of a parameter-dependent matrix $A(t)\in\mathbb{R}^{m\times n}$ with $m\geq n$ for a finite number of parameter values $t\in\mathbb{R}^{d}$ in some compact domain $D\subset\mathbb{R}^{d}$. This task has appeared in several applications, for example, in compression of a series of images [48], dynamical systems [36] and Gaussian random fields [38].
@@
where $X\in\mathbb{R}^{r\times n}$ is any row space approximator of $A$ and $Q_{X}\in\mathbb{R}^{n\times r}$ and $Q_{C}\in\mathbb{R}^{m\times r}$ are orthonormal matrices spanning the columns of $X^{T}$ and $C=A(:,J)$ respectively. [^2] This bound can also be stated in terms of a column approximator and a subset of the rows of $A$. The first two terms on the right-hand side of the inequality $\left\lVert Q_{C}(I,:)^{\dagger}\right\rVert_{2}$ and $\left\lVert Q_{X}(J,:)^{-1}\right\rVert_{2}$ mostly govern the accuracy of the CUR decomposition, because the row space approximator $X$ is arbitrary and can be e.g. chosen such that $\left\lVert A-AX^{\dagger}X\right\rVert_{F}$ is quasi-optimal. Therefore, it is important to get a good set of row and column indices that control the first two terms. There are many existing algorithms that lead to a good set of indices such as leverage score sampling [20, 42], DPP sampling [15], volume sampling [14, 16] and pivoting strategies [19, 21, 24, 27, 52, 53, 57], which all attempt to minimize the first two terms. Notably, there exists a set of row and column indices for which the CUR decomposition satisfies the following near-optimal guarantee [61]
@@
where $\llbracket{A}\rrbracket_{r}$ is the best rank-$r$ approximation to $A$. This means that little is lost by requiring our low-rank approximation to be in CUR form, as long as the indices are chosen appropriately. Polynomial-time algorithms for constructing such indices can be found in [14, 49]. The bound (2) can be improved by oversampling the rows, that is, obtain more row indices such that $I$ has more indices than the target rank. The topic of oversampling has been explored in [1, 23, 50, 51, 62]. Notably, [50] describes how to compute the CUR decomposition in a numerically stable way and show that oversampling can help improve the stability (in addition to accuracy) of the CUR decomposition.
@@
Now turning to parameter-dependent matrices $A(t)$, we can naively recompute the row and the column indices for each parameter value of interest. However, this approach can be inefficient and wasteful, as the computed indices for one parameter value are likely to contain information about nearby parameter values. The reason is because a matrix may only undergo slight deviations when the parameter value changes. More concretely, let $A(t)$ be a parameter-dependent matrix and $I$ and $J$ be a set of row and column indices respectively. Then by setting $X=V(t)^{T}$ where $V(t)$ is the $r$-dominant singular vectors of $A(t)$, (2) changes to
@@
(3) $\left\lVert A(t)-C(t)U(t)^{\dagger}R(t)\right\rVert_{F}\leq\left\lVert Q_{C(t)}(I,:)^{\dagger}\right\rVert_{2}\left\lVert Q_{V(t)}(J,:)^{-1}\right\rVert_{2}\left\lVert A(t)-\llbracket{A(t)}\rrbracket_{r}\right\rVert_{F}$
@@
where $\llbracket{A(t)}\rrbracket_{r}$ is the best rank-$r$ approximation to $A(t)$. Now the rightmost term in (3) is optimal for any parameter value $t$, so we compare the first two terms. Suppose we have two different parameter values, say $t_{1}$ and $t_{2}$, then if both $\left\lVert Q_{C(t_{1})}(I,:)^{\dagger}\right\rVert_{2}$ $\left\lVert Q_{V(t_{1})}(J,:)^{-1}\right\rVert_{2}$ and $\left\lVert Q_{C(t_{2})}(I,:)^{\dagger}\right\rVert_{2}\left\lVert Q_{V(t_{2})}(J,:)^{-1}\right\rVert_{2}$ remain small with the same set of indices $I$ and $J$, then we can use $I$ and $J$ to approximate both $A(t_{1})$ and $A(t_{2})$. It turns out that for nearby parameter values, the identical set of indices $I$ and $J$ often capture the important rows and columns of both $A(t_{1})$ and $A(t_{2})$, even in cases when $\left\lVert A(t_{1})-A(t_{2})\right\rVert_{2}$ is not small; see Section 4.333Since $\left\lVert A(t_{1})-A(t_{2})\right\rVert_{2}$ can be large in our setting, say $\mathcal{O}(1)$, perturbation bounds for the CUR decomposition in [30] is not enough to explain why the same indices can be reused for nearby parameter values. This motivates us to reuse the indices for various parameter values, potentially resulting in significant savings, often up to a factor $r$ in complexity where $r$ is the target rank. With this key idea in mind, the main goal of this paper is to devise a rank-adaptive certified [^4] We use the term certified to denote approximation methods that are designed to control the approximation error. algorithm that reuses the indices as much as possible until the error becomes too large, at which point we recompute the indices. To achieve this goal efficiently and reliably, we rely on a variety of efficient tools, many of which are borrowed from the randomized linear algebra literature such as randomized rank estimation [2, 45] and pivoting on a random sketch [19, 22, 57]; see Section 2. We assume throughout that for each parameter $t$, $A(t)$ has decaying singular values so that a low-rank approximation is possible. This is clearly a required assumption for low-rank approximation techniques to be effective.
@@
There have been several different approaches for finding a low-rank approximation to parameter-dependent matrices $A(t)$. We describe four different classes of methods. First, dynamical low-rank approximation is a differential-equation-based approach where the low-rank factors are obtained from projecting the time derivative of $A(t)$, $\dot{A}(t)$ onto the tangent space of the smooth manifold consisting of fixed-rank matrices; see [40]. A number of variations from the original dynamical low-rank approximation [36] have been proposed to deal with various issues such as the stiffness introduced by the curvature of the manifold [9, 35, 41] and rank-adaptivity [8, 31, 32, 33]. The complexity of this approach is typically $\mathcal{O}(r_{i}T_{A(t_{i})})$ for the parameter $t_{i}$, where $r_{i}$ is the target rank for $A(t_{i})$ and $T_{A(t_{i})}$ is the cost of matrix-vector product with $A(t_{i})$ or $A(t_{i})^{T}$. Secondly, Kressner and Lam [37] use randomized SVD [28] and generalized Nyström [46, 55], both based on random projections, to find low-rank approximations of parameter-dependent matrices $A(t)$. They focus on matrices that admit an affine linear decomposition with respect to $t$. Notably, they use the same random embedding for all parameter values rather than generating a new random embedding for each parameter value. The complexity is also $\mathcal{O}(r_{i}T_{A(t_{i})})$ for the parameter value $t_{i}$. Next, Donello et al. [17] use the CUR decomposition to find a low-rank approximation to parameter-dependent matrices. At each iteration, a new set of column and row indices for the CUR decomposition are computed by applying the discrete empirical interpolation method (DEIM) [10, 21, 52] to the singular vectors of the CUR decomposition from the previous iteration. While this approach benefits from not viewing the entire matrix, it may suffer from the same adversarial example as the fast algorithm, FastAdaCUR discussed in this paper; see Algorithm 6 and Section 4.4. They allow oversampling of column and/or row indices to improve the accuracy of the CUR decomposition and propose rank-adaptivity where the rank is either increased or decreased by $1$ if the smallest singular value of the current iterate is larger or smaller, respectively, than some threshold. Although their algorithm is based on the CUR decomposition, the low-rank factors in their approach are represented in the form of an SVD. This SVD is computed from the CUR decomposition at each iteration, given that orthonormal factors are necessary in DEIM. The complexity for this algorithm is $\mathcal{O}((m+n)r_{i}^{2})$ for the parameter value $t_{i}$. Lastly, in the special case when $A(t)$ is symmetric positive definite, a parametric variant of adaptive cross approximation has been used in [38].
@@
The contribution of this paper lies in the design of efficient rank-adaptive algorithms for a low-rank approximation of parameter-dependent matrices. The algorithms are based on the CUR decomposition of a matrix. We first demonstrate that the same set of row and column indices, or making slight modifications to the indices can yield effective low-rank approximations for nearby parameter values. This observation forms the basis of our efficient algorithms, which we describe in Section 3. We present two distinct algorithms: a rank-adaptive algorithm with error control, which we call AdaCUR (Adaptive CUR) and a faster rank-adaptive algorithm that sacrifices error control for speed, which we name FastAdaCUR (Fast Adaptive CUR); see Algorithms 5 and 6 respectively.
@@
FastAdaCUR, aside from the initial phase of computing the initial indices, runs linearly in $m$ and $n$ in the worse-case. The best-case complexity is $\mathcal{O}(r_{i}^{3})$ for the parameter value $t_{i}$, which makes the algorithm remarkably fast. While this algorithm is also rank-adaptive, it lacks rigorous error control since its priority is efficiency over accuracy. However, the algorithm has the advantage of needing only $\mathcal{O}(r)$ rows and columns for each parameter value, i.e., there is no need to view the full matrix. This feature is particularly attractive when the entries are expensive to compute. In experiments, we notice that the error may grow as we iterate through the parameter for difficult problems. Here, difficult problems refer to problems where $A(t)$ undergoes frequent large rank changes or one that changes rapidly such that the previous indices carry little to no information for the next. Nevertheless, we frequently observe that the algorithm performs well on easier problems; see Section 4.
@@
Throughout, we use $\left\lVert\cdot\right\rVert_{2}$ for the spectral norm or the vector-$\ell_{2}$ norm and $\left\lVert\cdot\right\rVert_{F}$ for the Frobenius norm. We use dagger to denote the pseudoinverse of a matrix and $\llbracket{A}\rrbracket_{r}$ to denote the best rank-$r$ approximation to $A$ in any unitarily invariant norm, i.e., the approximation derived from truncated SVD [34, § 7.4.9]. Unless specified otherwise, $\sigma_{i}(A)$ denotes the $i$th largest singular value of the matrix $A$. We use MATLAB style notation for matrices and vectors. For example, for the $k$th to $(k+j)$th columns of a matrix $A$ we write $A(:,k:k+j)$. Lastly, we use $|I|$ to denote the length of the vector or the cardinality of the set $I$, $I_{1}-I_{2}$ to be the set difference between $I_{1}$ and $I_{2}$, and define $[n]:=\{1,2,...,n\}$.

### v2

@@ Abstract
A low-rank approximation of a parameter-dependent matrix $A(t)$ is an important task in the computational sciences appearing for example in dynamical systems and compression of a series of images. In this work, we introduce AdaCUR, an efficient algorithm for computing a low-rank approximation of parameter-dependent matrices via CUR decompositions. The key idea for this algorithm is that for nearby parameter values, the column and row indices for the CUR decomposition can often be reused. AdaCUR is rank-adaptive, provides error control, and has complexity that compares favorably against existing methods. A faster algorithm which we call FastAdaCUR that prioritizes speed over accuracy is also given, which is rank-adaptive and has complexity which is at most linear in the number of rows or columns, but without error control.
@@ 1 Introduction
The task of finding a low-rank approximation to a matrix is ubiquitous in the computational sciences [56]. In large scale problems, low-rank approximation provides an efficient way to store and process the matrix. In this work, we study the low-rank approximation of a parameter-dependent matrix $A(t)\in\mathbb{R}^{m\times n}$ with $m\geq n$ for a finite number of parameter values $t\in\mathbb{R}$ in some compact domain $D\subset\mathbb{R}$. This task has appeared in several applications, for example, in compression of a series of images [48], dynamical systems [36] and Gaussian random fields [38].
@@
where $X\in\mathbb{R}^{r\times n}$ is any row space approximator of $A$ and $Q_{X}\in\mathbb{R}^{n\times r}$ and $Q_{C}\in\mathbb{R}^{m\times r}$ are orthonormal matrices spanning the columns of $X^{T}$ and $C=A(:,J)$ respectively. [^2] This bound can also be stated in terms of a column approximator and a subset of the rows of $A$. The first two terms on the right-hand side of the inequality $\left\lVert Q_{C}(I,:)^{\dagger}\right\rVert_{2}$ and $\left\lVert Q_{X}(J,:)^{-1}\right\rVert_{2}$ mostly govern the accuracy of the CUR decomposition, because the row space approximator $X$ is arbitrary and can be chosen such that $\left\lVert A-AX^{\dagger}X\right\rVert_{F}$ is quasi-optimal. Therefore, it is important to get a good set of row and column indices that control the first two terms. There are many existing algorithms that lead to a good set of indices such as leverage score sampling [20, 42], DPP sampling [15], volume sampling [14, 16] and pivoting strategies [19, 21, 24, 27, 52, 53, 57], which all attempt to minimize the first two terms. Notably, there exists a set of row and column indices for which the CUR decomposition satisfies the following near-optimal guarantee [61]
@@
where $\llbracket{A}\rrbracket_{r}$ is the best rank-$r$ approximation to $A$. This means that little is lost by requiring our low-rank approximation to be in CUR form, as long as the indices are chosen appropriately. Polynomial-time algorithms for constructing such indices can be found in [14, 49]. The bound (2) can be improved by oversampling the rows, that is, by obtaining more row indices such that $I$ has more indices than the target rank. The topic of oversampling has been explored in [1, 23, 50, 51, 62]. Notably, [50] describes how to compute the CUR decomposition in a numerically stable way and shows that oversampling can help improve the stability (in addition to accuracy) of the CUR decomposition.
Now, having defined what the CUR decomposition is, the objective of this work is the following: {objective} Let $A(t)$ be a parameter-dependent matrix. Then given a set of parameter values $t_{1},t_{2},...,t_{q}$, and a tolerance $\epsilon$, devise an algorithm that approximately achieves
(3) $\left\lVert A(t_{i})-\big{(}A(t_{i})(:,J_{i})\big{)}\left(A(t_{i})(I_{i},J_{i})^{\dagger}\right)\left(A(t_{i})(I_{i},:)\right)\right\rVert_{F}\leq\epsilon\left\lVert A(t_{i})\right\rVert_{F}$
for each $i$, where $I_{i}$ and $J_{i}$ are the row and column indices corresponding to $t_{i}$.
@@
For parameter-dependent matrices $A(t)$, we can naively recompute the row and the column indices for each parameter value of interest. However, this approach can be inefficient and wasteful, as the computed indices for one parameter value are likely to contain information about nearby parameter values. The reason is because a matrix may only undergo slight deviations when the parameter value changes. More concretely, let $A(t)$ be a parameter-dependent matrix and $I$ and $J$ be a set of row and column indices respectively. Then by setting $X=V_{r}(t)^{T}$ where $V_{r}(t)$ contains the $r$-dominant right singular vectors of $A(t)$, (2) changes to
(4) $\left\lVert A(t)-C(t)U(t)^{\dagger}R(t)\right\rVert_{F}\leq\left\lVert Q_{C(t)}(I,:)^{\dagger}\right\rVert_{2}\left\lVert Q_{V_{r}(t)}(J,:)^{-1}\right\rVert_{2}\left\lVert A(t)-\llbracket{A(t)}\rrbracket_{r}\right\rVert_{F}$
where $\llbracket{A(t)}\rrbracket_{r}$ is the best rank-$r$ approximation to $A(t)$. Now the rightmost term in (4) is optimal for any parameter value $t$, so we compare the first two terms. Suppose we have two different parameter values, say $t_{1}$ and $t_{2}$. Then, if both $\left\lVert Q_{C(t_{1})}(I,:)^{\dagger}\right\rVert_{2}$ $\left\lVert Q_{V_{r}(t_{1})}(J,:)^{-1}\right\rVert_{2}$ and $\left\lVert Q_{C(t_{2})}(I,:)^{\dagger}\right\rVert_{2}\left\lVert Q_{V_{r}(t_{2})}(J,:)^{-1}\right\rVert_{2}$ remain small with the same set of indices $I$ and $J$, we can use $I$ and $J$ to approximate both $A(t_{1})$ and $A(t_{2})$. [^3] In our work, we allow $\left\lVert A(t_{1})-A(t_{2})\right\rVert_{2}$ to be large, say $\mathcal{O}(1)$, and therefore the perturbation bounds for the CUR decomposition in [30] is not enough to explain why the same indices can be reused for nearby parameter values.
More specifically, define the following set of parameter-dependent matrices:
@@
(5) $\mathcal{M}_{IJ}^{(r)}(\delta):=\left\{A(t)\in\mathbb{R}^{m\times n}:\left\lVert Q_{A(t)(:,J)}(I,:)^{\dagger}\right\rVert_{2}\left\lVert Q_{V(t)}(J,:)^{-1}\right\rVert_{2}\leq\delta\text{ for all }t\in D\right\}$
for some domain $D\subset\mathbb{R}$ and sets of indices $I$ and $J$ with $|I|=|J|=r$. Then, for all $A(t)\in\mathcal{M}_{IJ}^{(r)}(\delta)$,
(6) $\left\lVert A(t)-C(t)U(t)^{\dagger}R(t)\right\rVert_{F}\leq\delta\left\lVert A(t)-\llbracket{A(t)}\rrbracket_{r}\right\rVert_{F}$
holds for all $t\in D$. Therefore if $A(t)\in\mathcal{M}_{IJ}^{(r)}(\delta)$ for sufficiently small $\delta\geq 1$ then the index sets $I$ and $J$ provide a good rank-$r$ CUR approximation to $A(t)$ for all $t\in D$. An example of a class of parameter-dependent matrices belonging to $\mathcal{M}_{IJ}^{(r)}(\delta)$ are incoherent rank-$r$ parameter-dependent matrices. In general, it is difficult to precisely determine whether a parameter-dependent matrix belongs to $\mathcal{M}_{IJ}^{(r)}(\delta)$ for some (practical) starting indices $I$ and $J$ or not, but intuitively, any parameter-dependent matrix that is sufficiently low-rank and has singular vectors that change gradually should belong to $\mathcal{M}_{IJ}^{(r)}(\delta)$ for some $I,J,\delta$ and $r$. We do not explore the technicalities of this further in this work.
@@
Nevertheless, in practice, for nearby parameter values, the identical set of indices $I$ and $J$ often capture the important rows and columns of both $A(t_{1})$ and $A(t_{2})$; see Sections 3 and 4. This motivates us to reuse the indices for various parameter values, potentially resulting in significant savings, often up to a factor $r$ in complexity where $r$ is the target rank. With this key idea in mind, the main goal of this paper is to devise a rank-adaptive certified [^4] We use the term certified to denote approximation methods that are designed to control the approximation error. algorithm that reuses the indices as much as possible until the error becomes too large, at which point we recompute the indices.
To achieve this goal efficiently and reliably, we rely on a variety of efficient tools, many of which are borrowed from the randomized linear algebra literature such as randomized rank estimation [2, 45] and pivoting on a random sketch [19, 22, 57]; see Section 2. We assume throughout that for each parameter $t$, $A(t)$ has decaying singular values so that a low-rank approximation is possible. This is clearly a required assumption for low-rank approximation techniques to be effective.
In the next section, we provide brief outlines of our algorithms, AdaCUR and FastAdaCUR, review existing methods for approximating parameter-dependent matrices, and summarize our contributions. This is followed by an overview of the various techniques from (randomized) numerical linear algebra that we use in our algorithms. The next two sections cover AdaCUR and FastAdaCUR in detail, as well as numerical experiments highlighting their performance. Finally, we conclude and discuss topics for further research.
1.1 Brief outline of AdaCUR and FastAdaCUR
Before introducing the two algorithms, AdaCUR and FastAdaCUR, in detail in Section 3, we first provide a brief outline of their core ideas.
AdaCUR is aimed at efficiently computing an accurate low-rank approximation of parameter-dependent matrices. A high-level overview of AdaCUR is given in Figure 1.
AdaCUR Input : Parameter-dependent matrix $A(t)\in\mathbb{R}^{m\times n}$, parameters $t_{1},...,t_{q}$, error tol. $\epsilon$. Output : CUR factors $C(t_{j}),U(t_{j}),R(t_{j})$ such that $A(t)\approx C(t_{j})U(t_{j})^{\dagger}R(t_{j})$. 1. Compute initial set of indices for $A(t_{1})$. For $j=2,3,...,q$,
Compute rel. error for $A(t_{j})$ using prev. indices Make minor modifications to the indices Recompute indices from scratch Rel. Err. $\leq\epsilon$Rel. Err. $>\epsilon$Rel. Err. $\leq\epsilon$Rel. Err. $>\epsilon$
Figure 1: An overview of AdaCUR
AdaCUR begins by computing an initial set of indices for $A(t_{1})$. For subsequent parameter values $t_{j}$, AdaCUR first computes the relative error for $A(t_{j})$ using the previous sets of indices. If the error is smaller than the tolerance $\epsilon$, the algorithm retains the previous sets of indices and proceeds to the next parameter value. If the error exceeds the tolerance, minor modifications are made to the indices, and the algorithm checks whether the tolerance is satisfied. If the modified indices meet the tolerance, they are used; otherwise, the indices are recomputed from scratch. The precise details of AdaCUR are outlined in Section 3.2.
FastAdaCUR is aimed at achieving even greater speed at the cost of possible reduction in accuracy. Instead of computing the error which can be expensive, FastAdaCUR keeps a small set of extra indices that helps with rank adaptivity. A high-level overview of FastAdaCUR is given in Figure 2.
FastAdaCUR Input : Parameter-dependent matrix $A(t)\in\mathbb{R}^{m\times n}$, parameters $t_{1},...,t_{q}$, rank tol. $\epsilon$ Output : CUR factors $C(t_{j}),U(t_{j}),R(t_{j})$ such that $A(t)\approx C(t_{j})U(t_{j})^{\dagger}R(t_{j})$. 1. Compute initial set of indices $I,J$ and a small set of extra indices $I_{s},J_{s}$ for $A(t_{1})$. For $j=2,3,...,q$,
Compute $\epsilon$-rank of $A(I\cup I_{s},J\cup J_{s})$ Add indices to $I,J$ from $I_{s},J_{s}$ and add more indices to $I_{s},J_{s}$ Remove indices from $I,J$ $\epsilon$-rank decr.$\epsilon$-rank incr.
Figure 2: An overview of FastAdaCUR
FastAdaCUR begins by computing an initial sets of indices $I,J$ and a small set of additional indices $I_{s},J_{s}$ for $A(t_{1})$. For subsequent parameter values $t_{j}$, FastAdaCUR calculates the $\epsilon$-rank of the core matrix $A(I\cup I_{s},J\cup J_{s})$ using the previous sets of indices to determine whether the $\epsilon$-rank exceeds the size of the index sets $I$ and $J$. Importantly, for FastAdaCUR, the entire matrix is not accessed beyond the first parameter value, which makes the algorithm efficient. If the $\epsilon$-rank has increased, indices are added to $I$ and $J$ from $I_{s}$ and $J_{s}$. Conversely, if the $\epsilon$-rank has decreased, indices are removed from $I$ and $J$. After this adjustment, the algorithm proceeds to the next parameter value. The details of FastAdaCUR are provided in Section 3.3. Unlike AdaCUR, FastAdaCUR does not have error control mechanism, making it vulnerable to adversarial examples; see Section 4.4.
@@ 1.1 Brief outline of AdaCUR and FastAdaCUR
There have been several different approaches for finding a low-rank approximation to parameter-dependent matrices $A(t)$. We describe four different classes of methods. First, dynamical low-rank approximation is a differential-equation-based approach where the low-rank factors are obtained from projecting the time derivative of $A(t)$, $\dot{A}(t)$, onto the tangent space of the smooth manifold consisting of fixed-rank matrices; see [40]. A number of variations from the original dynamical low-rank approximation [36] have been proposed to deal with various issues such as the stiffness introduced by the curvature of the manifold [9, 35, 41] and rank-adaptivity [8, 31, 32, 33]. The complexity of this approach is typically $\mathcal{O}(r_{i}T_{A(t_{i})})$ for the parameter $t_{i}$, where $r_{i}$ is the target rank for $A(t_{i})$ and $T_{A(t_{i})}$ is the cost of matrix-vector product with $A(t_{i})$ or $A(t_{i})^{T}$. Secondly, Kressner and Lam [37] use randomized SVD [28] and generalized Nyström [46, 55], both based on random projections, to find low-rank approximations of parameter-dependent matrices $A(t)$. They focus on matrices that admit an affine linear decomposition with respect to $t$. Notably, they use the same random embedding for all parameter values rather than generating a new random embedding for each parameter value. The complexity is also $\mathcal{O}(r_{i}T_{A(t_{i})})$ for the parameter value $t_{i}$. Next, Donello et al. [17] use the CUR decomposition to find a low-rank approximation to parameter-dependent matrices. At each iteration, a new set of column and row indices for the CUR decomposition are computed by applying the discrete empirical interpolation method (DEIM) [10, 21, 52] to the singular vectors of the CUR decomposition from the previous iteration. While this approach benefits from not viewing the entire matrix, it may suffer from the same adversarial example as the fast algorithm, FastAdaCUR discussed in this paper; see Algorithm 7 and Section 4.4. They allow oversampling of column and/or row indices to improve the accuracy of the CUR decomposition and propose rank-adaptivity where the rank is either increased or decreased by $1$ if the smallest singular value of the current iterate is larger or smaller, respectively, than some threshold. Although their algorithm is based on the CUR decomposition, the low-rank factors in their approach are represented in the form of an SVD. This SVD is computed from the CUR decomposition at each iteration, given that orthonormal factors are necessary in DEIM. The complexity for this algorithm is $\mathcal{O}((m+n)r_{i}^{2})$ for the parameter value $t_{i}$. Lastly, in the special case when $A(t)$ is symmetric positive definite, a parametric variant of adaptive cross approximation has been used in [38].
@@
The contribution of this paper lies in the design of efficient rank-adaptive algorithms for low-rank approximations of parameter-dependent matrices. The algorithms are based on the CUR decomposition of a matrix. We first demonstrate that the same set of row and column indices, or making slight modifications to the indices, can yield effective low-rank approximations for nearby parameter values. This observation forms the basis of our efficient algorithms, which we describe in Section 3. We present two distinct algorithms: a rank-adaptive algorithm with error control, which we call AdaCUR (Adaptive CUR), and a faster rank-adaptive algorithm that sacrifices error control for speed, which we name FastAdaCUR (Fast Adaptive CUR); see Algorithms 6 and 7 respectively.
@@
FastAdaCUR, aside from the initial phase of computing the initial indices, runs linearly in $m$ and $n$ in the worse-case. The best-case complexity is $\mathcal{O}(r_{i}^{3})$ for the parameter value $t_{i}$, which makes the algorithm remarkably fast. While this algorithm is also rank-adaptive, it lacks rigorous error control since its priority is efficiency over accuracy. However, the algorithm has the advantage of needing only $\mathcal{O}(r)$ rows and columns for each parameter value; i.e., there is no need to view the full matrix. This feature is particularly attractive when the entries are expensive to compute. In experiments, we notice that the error may grow as we iterate through the parameter for difficult problems. Here, difficult problems refer to problems where $A(t)$ undergoes frequent large rank changes or one that changes rapidly such that the previous indices carry little to no information for the next. Nevertheless, we frequently observe that the algorithm performs well on easier problems; see Section 4.
@@
Throughout, we use $\left\lVert\cdot\right\rVert_{2}$ for the spectral norm or the vector-$\ell_{2}$ norm and $\left\lVert\cdot\right\rVert_{F}$ for the Frobenius norm. We use dagger to denote the pseudoinverse of a matrix and $\llbracket{A}\rrbracket_{r}$ to denote the best rank-$r$ approximation to $A$ in any unitarily invariant norm, i.e., the approximation derived from truncated SVD [34, § 7.4.9]. We use $a_{1}\lesssim a_{2}$ to denote that $a_{1}\leq Ca_{2}$ for some constant $C>0$. Unless specified otherwise, $\sigma_{i}(A)$ denotes the $i$th largest singular value of the matrix $A$. We use MATLAB style notation for matrices and vectors. For example, for the $k$th to $(k+j)$th columns of a matrix $A$ we write $A(:,k:k+j)$. Lastly, we use $|I|$ to denote the length of the vector or the cardinality of the set $I$, $I_{1}-I_{2}$ to be the set difference between $I_{1}$ and $I_{2}$, and define $[n]:=\{1,2,...,n\}$.